OpenStudy (anonymous):

how to graph f(x) = 2/3x - 1/2 Please help me. Really need some help here.

OpenStudy (abmon98):

Set x=0 and y=0 to find both y-intercept and x-intercept y=2/3(0)-1/2 y=-1/2 0=2/3x-1/2 2/3x=1/2 x=1/2*3/2=3/4 

OpenStudy (anonymous):

But can you use the formula m=rise/run here?

OpenStudy (mathmath333):

\[y=slope (rise/run)x +b\]

OpenStudy (anonymous):

So will the graph still be the same if I use the formula you have given and when I solve by substituting x and y by O in finding the x and y intercept?

OpenStudy (mathmath333):

yes absolutely

OpenStudy (anonymous):

Ahhh.. Okay, thanks :)